# There is No Silver Bullet: Benchmarking Methods in Predictive Combinatorial Optimization

### v4

@@ Abstract
Predictive combinatorial optimization, where the parameters of combinatorial optimization (CO) are unknown at the decision-making time, is the precise modeling of many real-world applications, including energy cost-aware scheduling and budget allocation on advertising. Tackling such a problem usually involves a prediction model and a CO solver. These two modules are integrated into the predictive CO pipeline following two design principles: “Predict-then-Optimize (PtO)”, which learns predictions by supervised training and subsequently solves CO using predicted coefficients, while the other, named “Predict-and-Optimize (PnO)”, directly optimizes towards the ultimate decision quality and claims to yield better decisions than traditional PtO approaches. However, there lacks a systematic benchmark of both approaches, including the specific design choices at the module level, as well as an evaluation dataset that covers representative real-world scenarios. To this end, we develop a modular framework to benchmark 11 existing PtO/PnO methods on 8 problems, including a new industrial dataset for combinatorial advertising that will be released. Our study shows that PnO approaches are better than PtO on 7 out of 8 benchmarks, but there is no silver bullet found for the specific design choices of PnO. A comprehensive categorization of current approaches and integration of typical scenarios are provided under a unified benchmark. Therefore, this paper could serve as a comprehensive benchmark for future PnO approach development and also offer fast prototyping for application-focused development.
@@ 1 Introduction
Hence, there are important and practical needs to jointly consider both prediction and decision-making for predictive optimization (especially CO) problems [18, 36, 10]. A basic solution for predictive CO is “predict-then-optimize” [7](“PtO”, or called the two-stage approach), which first predicts coefficients of the optimization task through a predictive model trained under the supervision of ground truth coefficients and subsequently utilizes off-the-shelf solvers to obtain solutions. Intuitively, it is expected that a higher prediction accuracy would result in better decision quality. Nonetheless, as evidenced by Figure 5(c~d), there often exists a gap between prediction objectives and ultimate decision goals, leading to suboptimal decisions by PtO models. Therefore, as shown in Fig 1(b), a recent series of studies [36, 55] have proposed the new “predict-and-optimize” paradigm (“PnO”, or decision-focused learning [57, 34, 50]) that learns to cater for the ultimate decision objectives and exhibits notable improvement over PtO on several optimization tasks.
@@
In this work, we systematically review existing methodologies and establish a benchmark to align each model. Based on empirical results, we then conduct a series of in-depth experiments to explore key factors for PnO model design, and find several useful tips for deployment of PnO models. Based on these findings, we provide a set of recommendations for practitioners on model selection, considering the applicability and practical performance. We anticipate that our open-source benchmark  [^1] https://anonymous.4open.science/r/PTO-0B23/ and new dataset will gain increased attention, thereby fostering advancements in both research and practical applications. Our contributions and conclusions are summarized as follows:
@@
$\bullet$   Our benchmark results demonstrate PnO is better than PtO on 7 out of 8 predictive CO problems. However, no silver bullet is found for specific design choices, suggesting necessity of trial-and-error for various scenarios. Therefore, our comprehensive and modular-based benchmark covering mainstream PnOs could further help quick prototyping in application-focused development.
@@
Table 1: Categorization of representative methods, where “Predictive” belongs to “PtO” while others are “PnO”s. BB, ID, CPLayer is short for Blackbox, Identity, and cvxpy layer. “✓✗” means it supports multiple versions. Suppose $N$ is the number of decision variables (or constraints), and $K$ is the number of extra optimization samples. $T_{LODL}$ denotes backward time following  [50].
@@ 2 Problem Formulation
$\min_{\mathbf{z}\in\mathcal{Z}}\quad f(\mathbf{z},\mathbf{y},\bm{\theta})~{}s.t.~{}\operatorname{Constr}(\mathbf{z},\bm{\theta})~{},$ (1)
where $f(\mathbf{z},\mathbf{y},\bm{\theta})$ denotes the known and closed-formed optimization objective (abbreviated as $f(\mathbf{z},\mathbf{y})$ below) with discrete variable $\mathbf{z}\in\mathcal{Z}$, and $\mathbf{y}$ is collection of unknown optimization coefficients, $\bm{\theta}$ are optimization parameters that are known and fixed, and $\operatorname{Constr}(\mathbf{z},\bm{\theta})$ are the constraints. We assume that the coefficients of the constraints are known in line with the majority of literature [57, 50, 34]. We denote one solver call for a problem instance with coefficient $\mathbf{y}$ as $\mathbf{z}(\mathbf{y})$, and $\mathbf{z}^{\star}(\mathbf{y})$ as the optimal solution. The solvers can be commercial solvers (e.g. Gurobi [25]), open-sourced solvers (e.g. cvxpy [15]), or neural solvers (e.g. submodular [30]).
@@
The evaluation of the PtO could be critical. In many circumstances where the coefficients of the test set are available, regret [36, 60, 24, 34] is used to evaluate decision quality. Let the decision quality [50, 21] of solution $\hat{\mathbf{z}}$ be the objective under the ground-truth coefficient $\mathbf{y}$:
@@
Then the regret could be obtained by the difference of the decision quality with solutions under the estimated coefficient($\mathbf{z}^{\star}(\mathbf{y}),$ and ground-truth coefficients ($\mathbf{z}^{\star}(\hat{\mathbf{y}})$):
@@
However, in many cases (e.g. combinatorial advertising problem), the ground truth coefficients ${\mathbf{y}}$ are not readily available or even impossible to obtain, so the evaluation of PtO becomes challenging. Online A/B testing may be one approach; however, it is associated with higher costs and risks in practice. Consequently, we propose using uplift, an offline controlled variable metric common in causal inference tasks, as an alternative evaluation elaborated in Appendix B.7.3.
@@ 3 Categorization of Existing Methods
As briefly introduced earlier, the major challenges of PnO lie in two aspects: (1) unavailable/non-informative gradients and (2) discrete variables. These two challenges both block the gradient backpropagation so that the prediction model cannot be updated by the final objective. As shown in Table 1, we categorize existing works into four categories according to how the optimization is solved and how the gradient is obtained, where the gradient flow of these approaches is shown in Figure 2. We consider the PtO training as a baseline method while concurrently noting that the first challenge can be readily addressed if the solver is differentiable, such as neural solvers for top-k [58] or Sinkhorn algorithm [51] used in neural graph matching [56]. The technical details are in Appendix A.3.
Figure 2: Gradient flow of existing PnO methods, where $f$, $\hat{f}$ and $\tilde{f}$ denote the original objective, learned objective by surrogate model, and the continuous relaxation of the original objective, respectively. The path of back propagation of the vanilla two-stage is ①, while discrete/continuous categories use ①②, statistical one uses  ④, and surrogate one uses  ①③.
@@ 3.1 The Discrete Category
As discussed in previous works [44, 6, 40], the gradient of decision variable $\mathbf{z}$ with respect to coefficients $\mathbf{y}$ revealed as piecewise constant functions (almost zero everywhere, and undefined otherwise). To address this issue, the “discrete” category solves the optimization with the original discrete solver in the forward pass, while in the backward pass, the gradients are estimated via designed interpolation functions, such as linear interpolation done by Blackbox [44]. Representative models include DFL [50], Blackbox [44], Perturb [6] and I-MLE [40], Identity [47], etc.
@@ 3.3 The Statistical Category
The “statistical” category [39, 34] designs loss functions for PnO considering the statistical relationships between multiple solutions. The aim of designed losses is to make coefficient predictions so that the decision quality (defined in Eq. (4)) of optimal solutions is superior to suboptimal solutions. Representative models include NCE [39] and LTR [34] (with 3 versions: pointwise, pairwise, and listwise-LTR).
The statistical category is usually agnostic to the problem type and solver. However, it requires a solution cache before the training, which creates additional overhead prior to training. It is also not readily applicable to problem instances of varying variable sizes for solution cache in implementation.
@@ 3.4 The Surrogate Category
Similar to the statistical category, the surrogate category easily adapts to all problem and solver types. Furthermore, in cases where solving is slow or assessing solution quality is expensive, using the surrogate expedites the PnO training process. Nevertheless, they do necessitate a set of optimization instances and corresponding solutions to learn objective functions as initialization. Additionally, whether they can learn complex nonlinear functions remains open.
@@ 3.5 Takeaways for practitioners’ PnO model choice
PnO model choice guideline As shown in Fig 4, the preferable model depends on requirement for labeled prediction datasets, extra optimization samples, and overall performance (both efficacy and efficiency). Firstly, the SPO [36] method is preferable for most linear or near-linear problems (Knapsack, energy-cost aware scheduling, budget allocation problem, etc.) as the first try, as long as there are labeled datasets for prediction tasks. SPO requires fewer labeled datasets and is generally faster regarding additional solving, backward pass complexity, etc. Secondly, the LTR [34] method is a good choice for non-linear tasks (TopK, bipartite matching, Portfolio optimization) as long as there are extra optimization samples and a solution cache, and LODL [50] for better performance if optimization solutions are easy to get (since it requires abundant samples and solutions). Finally, the discrete category (Blackbox [44]/Identity [47]) is the last choice for linear/convex problems if there are no labeled datasets for prediction tasks, such as combinatorial advertising.
4 Benchmarks and Rethinkings
@@ 4.1 Modular Framework Design
As shown in Figure 3, we implement the benchmark as a modular framework where each scenario is composed of problems, predictors, solvers, losses, and evaluations. This enables a comprehensive and convenient evaluation of current methods. Although we have not exhaustively addressed all CO problems, we have selected representative ones from real-world applications and introduced an authentic industrial dataset to contribute to the field’s development.
@@ 4.2 Problem Descriptions
In this section, we briefly introduce each task with the datasets in Table 2, where the background, practical scenarios, detailed “prediction” and “optimization” formulations are left in Appendix B.
@@ 4.2.1 Benchmark on publicly available datasets
Budget Allocation (BA) BA has applications when agents intend to disseminate information [57, 50] across multiple websites, such as Facebook, etc., to reach a wider audience within a budget constraint, but the probability of each website reaching users is unknown.
@@ 4.2.2 New dataset of combinatorial advertising for inclusive finance
Dataset: The advertisement takes place on a mobile application (APP) as a combination of various channels, including in-app notifications, text messages, telephones, etc. The outcome of each combination for each user is unknown, and each channel is associated with its respective cost. The data will be publicly accessible after publication, and the processing and use terms are elaborated in Appendix B.7.
Prediction: Given user $i$’s feature $\mathbf{x}^{i}$ (encrypted and processed personal features, app activity records, etc.), predict the user’s conversion rate $\mathbf{y}^{ij}$ to the $j$-th advertising combination.
@@
This problem differs from the above BA problem as CA is based on personalized advertising while BA is not. CA is also more challenging than the multiple treatment setting [63] due to the exponential combination space of channels, where the latter refers to multiple levels of treatments of the same channel.
@@ 4.3 Experimental Setup
For the predictive model, we use a two-layer MLP with 32 hidden units in each layer. Unless otherwise specified, we extract 20% from the training data set for validation. During training, we adopt the Adam Optimizer [31] and search by grid the learning rate in $\{5e-2,1e-2,5e-3,1e-3\}$. In each run, each model is trained for 300 epochs and the training stops if no better regret is achieved for 50 epochs on the validation set. The epoch that reaches the lowest regret (or highest uplift) in validation is selected for testing. We elaborate on more details of the model design in Appendix A.
@@ 4.4 Benchmark Results
We list the benchmarks, including relative regret (with respect to optimal objective) and average runtime for training and test in Table 3, and hyper-parameter sensitivity in Appendix C.3.
From the perspective of each method, methods in the discrete category do not perform as well as others, especially on non-linear optimization tasks (budget allocation, topk, portfolio, etc.). This is probably because the gradient interpolations are only designed for linear objectives and gradient estimations in the back-propagation may not be accurate in complex problems. The continuous-category method SPO performs well on linear problems with relatively short training time. We run CPLayer on the linear optimization problems and non-linear programs are omitted. For the statistical category, listwise-LTR achieves better results than others in most problems, probably because it captures the global relationships among multiple solutions. The surrogate method LODL achieves satisfactory results on half of the benchmarks. More analysis is left to Appendix C.1.
@@
The result of the combinatorial advertising on discrete category is shown in Table 4, where other categories of methods that cannot run on the fly do not apply to this problem directly because they require additional coefficient labels/optimization samples that cannot be satisfied in this problem. We conclude applicability in Sec. 3.5 in the following. The PnO training significantly improves the uplift metric (using the discrete category). The identity model achieves the best result. The DFL and Blackbox approach also performs better than the two-stage approach, though with higher training time. Though the end-to-end PnO training takes higher training time, they do not incur additional runtime during the test stage.
@@
Based on the above categorizations and benchmark results, we find that PnO methods lack a universally accepted standard, and in practical applications, each method has its own advantages and disadvantages. Therefore, there is an urgent need for better approaches with respect to applicability, efficiency and performance. Hence, addressing the imperative challenge of attaining a more universal, stable, and efficient training model is essential to fully realize the potential of “PnO” in the real world.
@@
4.5 Key factors for future PnO methodology design

### v5

# Benchmarking PtO and PnO Methods in the Predictive Combinatorial Optimization Regime

@@ Abstract
Predictive combinatorial optimization, where the parameters of combinatorial optimization (CO) are unknown at the decision-making time, is the precise modeling of many real-world applications, including energy cost-aware scheduling and budget allocation on advertising. Tackling such a problem usually involves a prediction model and a CO solver. These two modules are integrated into the predictive CO pipeline following two design principles: “Predict-then-Optimize (PtO)”, which learns predictions by supervised training and subsequently solves CO using predicted coefficients, while the other, named “Predict-and-Optimize (PnO)”, directly optimizes towards the ultimate decision quality and claims to yield better decisions than traditional PtO approaches. However, there lacks a systematic benchmark of both approaches, including the specific design choices at the module level, as well as an evaluation dataset that covers representative real-world scenarios. To this end, we develop a modular framework to benchmark 11 existing PtO/PnO methods on 8 problems, including a new industrial dataset for combinatorial advertising that will be released. Our study shows that PnO approaches are better than PtO on 7 out of 8 benchmarks, but there is no silver bullet found for the specific design choices of PnO. A comprehensive categorization of current approaches and integration of typical scenarios are provided under a unified benchmark. Therefore, this paper could serve as a comprehensive benchmark for future PnO approach development and also offer fast prototyping for application-focused development. The code is available at https://github.com/Thinklab-SJTU/PredictiveCO-Benchmark.
@@ 1 Introduction
Hence, there are important and practical needs to jointly consider both prediction and decision-making for predictive optimization (especially CO) problems [20, 46, 9]. A basic solution for predictive CO is “predict-then-optimize” [7] (PtO, or called the two-stage approach), which first predicts coefficients of the optimization task through a predictive model trained under the supervision of ground truth coefficients and subsequently utilizes off-the-shelf solvers to obtain solutions. Intuitively, it is expected that a higher prediction accuracy would result in better decision quality. Nonetheless, as also evidenced by Figure 5(c~d), there often exists a gap between prediction objectives and ultimate decision goals, leading to suboptimal decisions by PtO models. Therefore, as shown in Fig 1(b), a recent series of studies [46, 68] have proposed the new “predict-and-optimize” paradigm (PnO, or decision-focused learning [72, 44, 60] that learns to directly cater for the ultimate decision objectives. In cases where the quality of the final decision is paramount, joint prediction-and-optimization approaches may be more direct and beneficial. Though this can sometimes come at the expense of prediction accuracy (as is shown in Fig 5(c)), they exhibit notable improvement over PtO on several optimization tasks.
@@
In this work, we systematically review existing methodologies and establish a benchmark to align each model. Based on empirical results, we then conduct a series of in-depth experiments to explore key factors for PnO model design, and find several useful tips for deployment of PnO models. Based on these findings, we provide a set of recommendations for practitioners on model selection when considering the applicability and practical performance. We anticipate that our open-source benchmark and new dataset will gain increased attention, thereby fostering advancements in both research and practical applications. Our contributions and conclusions are summarized as follows:
@@
$\bullet$    Our benchmark results demonstrate PnO is better than PtO on 7 out of 8 predictive CO problems. However, no silver bullet is found for specific design choices, suggesting the necessity of trial-and-error for various scenarios. Therefore, our comprehensive and modular-based benchmark covering mainstream PnOs could further help quick prototyping in application-focused development.
@@
Table 1: Categorization of representative methods, where “Predictive” belongs to “PtO” while others are “PnO”s. BB, ID, CPLayer is short for Blackbox, Identity, and cvxpy layer. “✓✗” means it supports the characteristic in some models among many. Suppose $N$ is the decision variable size, and $K$ is the number of extra optimization samples. $T_{LODL}$ denotes backward time following  [60].
@@ 2 Problem Formulation
$\max_{\mathbf{z}\in\mathcal{Z}}\quad f(\mathbf{z},\mathbf{y},\bm{\theta})~{}s.t.~{}\mathbf{z}\in\operatorname{Constr}(\bm{\theta})~{},$ (1)
where $f(\mathbf{z},\mathbf{y},\bm{\theta})$ denotes the known and closed-formed optimization objective (abbreviated as $f(\mathbf{z},\mathbf{y})$ below) with discrete variable $\mathbf{z}\in\mathcal{Z}$, and $\mathbf{y}$ is collection of unknown optimization coefficients, $\bm{\theta}$ are optimization parameters that are known and fixed, and decision variables $\mathbf{z}$ obey the constraints $\operatorname{Constr}(\bm{\theta})$. We assume that the coefficients of the constraints are known in line with the majority of literature [72, 60, 44]. We denote the solution obtained through one solver call for a problem instance with coefficient $\mathbf{y}$ as $\mathbf{z}(\mathbf{y})$, and $\mathbf{z}^{\star}(\mathbf{y})$ as the optimal solution. The solvers can be commercial solvers (e.g. Gurobi [28]), open-sourced solvers (e.g. cvxpy [16]), or neural solvers (e.g. submodular [37]).
@@
The evaluation of the PtO could be critical. In many circumstances where the real coefficients of the test set are available, regret [46, 75, 26, 44] is used to evaluate decision quality. Let the decision quality [60, 23] of solution $\hat{\mathbf{z}}$ be the objective under the ground-truth coefficient $\mathbf{y}$:
@@
Then the regret could be obtained by the difference of the decision quality with solutions under the estimated coefficient($\mathbf{z}^{\star}(\mathbf{y}),$) and ground-truth coefficients ($\mathbf{z}^{\star}(\hat{\mathbf{y}})$):
@@
However, in many cases (e.g. combinatorial advertising problem), the ground truth coefficients ${\mathbf{y}}$ are not readily available or even impossible to obtain, so the evaluation of PtO becomes challenging. Online A/B testing may be one approach; however, it is associated with higher costs and risks in practice. Consequently, we propose using uplift, an offline controlled variable metric common in causal inference tasks, as an alternative evaluation elaborated in Appendix C.7.3.
@@ 3 Categorization of Existing Methods
As briefly introduced earlier, the major challenges of PnO lie in two aspects: (1) unavailable/non-informative gradients and (2) discrete variables. These two challenges both block the gradient backpropagation so that the prediction model cannot be updated by the final objective. Table 1 categorizes existing works into four groups by how the optimization is solved and how the gradient is obtained, where the gradient flow of these approaches is shown in Figure 2. We consider the PtO training as a baseline method while concurrently noting that the first challenge can be readily addressed if the solver is differentiable, such as neural solvers for top-k [73] or Sinkhorn algorithm [62] used in neural graph matching [70]. The technical details are given in Appendix B.3.
Figure 2: Gradient flow of existing PnO methods, where $f$, $\hat{f}$ and $\tilde{f}$ denote the original objective, learned objective by surrogate model, and the continuous relaxation of the original objective, respectively. The path of back propagation of the vanilla two-stage is ①, while discrete/continuous categories use ①②, statistical one uses  ④, and surrogate one uses  ①③. $l_{surro}$ represents the surrogate losses to measure how the surrogate imitates the original optimization objective, while $l_{rank}$ refers to the designed loss that encodes solution ranking—e.g., ensuring that the optimal solution is assigned a lower loss than suboptimal ones.
@@ 3.1 The Discrete Category
As discussed in [54, 6, 50], the gradient of decision variable $\mathbf{z}$ with respect to coefficients $\mathbf{y}$ revealed as piecewise constant functions (almost zero everywhere, and undefined otherwise). To address this issue, the “discrete” category solves the optimization with the original discrete solver in the forward pass, while in the backward pass, the gradients are estimated via designed interpolation functions, such as linear interpolation done by Blackbox [54]. Representative models include DFL [60], Blackbox [54], Perturb [6] and I-MLE [50], Identity [58], etc.
@@ 3.3 The Statistical Category
The “statistical” category [49, 44] designs loss for PnO considering the statistical relation between multiple solutions. The aim is to make coefficient predictions so that the decision quality (defined in Eq. (4)) of optimal solutions is superior to suboptimal solutions. Representative models include NCE [49] and LTR [44] (with 3 versions: pointwise-LTR, pairwise-LTR, and listwise-LTR).
The statistical category is usually agnostic to the problem type and solver. However, a solution cache of optimization samples is required before the training, which creates additional overhead. It is also not readily applicable to problem instances of varying variable sizes for solution cache in implementation.
@@ 3.4 The Surrogate Category
Similar to the statistical category, the surrogate category easily adapts to all problem and solver types. Furthermore, in cases where solving is slow or assessing solution quality is expensive, using the surrogate expedites the PnO training process. Nevertheless, they do necessitate an extra set of optimization instances and corresponding solutions to learn objective functions as initialization. Additionally, whether and how well they can learn complex nonlinear functions remains open.
@@ 3.5 Takeaways for Practitioners’ PnO Model Choice
PnO model choice guideline As shown in Fig 4, we consider several key factors for the preferable model recommendation: requirement for labeled prediction datasets, extra optimization samples, and overall performance (in both efficacy and efficiency). Firstly, the SPO [46] method is preferable for most linear or near-linear problems (Knapsack, energy-cost aware scheduling, budget allocation problem, etc.) as the first try, as long as there are labeled datasets for prediction tasks. SPO requires fewer labeled datasets and is generally faster regarding additional solving, backward pass complexity, etc. Secondly, the LTR [44] method is a good choice for non-linear tasks (TopK, bipartite matching, Portfolio optimization) as long as there are extra optimization samples at hand (or corresponding solution cache), and LODL [60] for better performance if optimization solutions are easy to get (since it requires abundant samples and solutions). Finally, the discrete category (Blackbox [54]/Identity [58]) is the last choice for linear/convex problems if there are no labeled datasets for prediction tasks, such as combinatorial advertising.
4 Benchmarks and Remarks
@@ 4.1 Modular Framework Design
As shown in Figure 3, we implement the benchmark as a modular framework where each scenario is composed of problems, predictors, solvers, losses, and evaluations. This enables a comprehensive and convenient evaluation of current methods. Although we have not exhaustively addressed all CO problems, we have selected representative ones from real-world applications and introduced a real-world industrial dataset to contribute to the field’s development.
@@ 4.2 Problem Descriptions
In this section, we briefly introduce each task with the datasets in Table 2, where the background, practical scenarios, detailed “prediction” and “optimization” formulations are left in Appendix C.
@@ 4.2.1 Benchmark on publicly available datasets
Budget Allocation (BA) BA has applications when agents intend to disseminate information [72, 60] across multiple websites to reach a wider audience within a budget constraint, but the probability of each website reaching users is unknown.
@@ 4.2.2 New dataset of combinatorial advertising for inclusive finance
Dataset: The advertisement takes place on a mobile application (APP) as a combination of various channels, including in-app notifications, text messages, telephones, etc. Historical data contains whether a user converted after being exposed to a specific marketing combination in the past; however, labels for other combinations remain unknown. The data can be accessed through the website [^1] https://opendata.xinye.com, and the processing and use terms are in Appendix C.7.
Prediction: Given user $i$’s feature $\mathbf{x}^{i}$ (encrypted and processed personal features, app activity records, etc.), predict the user’s conversion rate $\mathbf{y}^{ij}$ to the $j$-th combination of advertising channels.
@@
It differs from the above BA problem as CA is based on personalized advertising while BA is not. CA is also more challenging than the multiple treatment setting [79] due to the exponential combination space of channels, where the latter refers to multiple levels of treatments of the same channel.
@@ 4.3 Experimental Setup
[…]
@@ 4.4 Benchmark Results
We list the benchmarks, including relative regret (with respect to optimal objective) and average runtime for training and test in Table 3, and hyper-parameter sensitivity in Appendix D.3.
From the perspective of each method, methods in the discrete category do not perform as well as others, especially on non-linear optimization tasks (budget allocation, topk, portfolio, etc.). This is probably because the gradient interpolations are only designed for linear objectives and gradient estimations in the back-propagation may not be accurate in complex problems. The continuous-category method SPO performs well on linear problems with relatively short training time. We run CPLayer on the linear optimization problems and non-linear programs are omitted. For the statistical category, listwise-LTR achieves better results than others in most problems, probably because it captures the global relationships among multiple solutions. The surrogate method LODL achieves satisfactory results on half of the benchmarks. More analysis is left to Appendix D.1.
@@
The result of the combinatorial advertising on discrete category is shown in Table 4, where other categories of methods that cannot run on-the-fly do not apply to this problem directly because they require additional coefficient labels/optimization samples that cannot be satisfied in this problem. We conclude applicability in Sec. 3.5 in the following. The PnO training significantly improves the uplift metric (using the discrete category). The identity model achieves the best result. The DFL and Blackbox approach also performs better than the two-stage approach, though with higher training time. Though the end-to-end PnO training takes higher training time, they do not incur additional runtime during the test stage.
@@
The above categorizations and benchmark suggest that PnO methods lack a universally accepted standard, and in practical applications, each method has its own advantages and disadvantages. Therefore, there is an urgent need for better approaches with respect to applicability, efficiency and performance. Hence, addressing the imperative challenge of attaining a more universal, stable, and efficient training model is essential to fully realize the potential of “PnO” in the real world.
@@
4.5 Key Factors for PnO Methodology Design